OpenStudy (anonymous):

write an equation in standard form with slope 1/4 and y-intercept -3

OpenStudy (anonymous):

y=1/4x - 3

OpenStudy (anonymous):

First put it in Y-Intercept - y=1/4x-3 then turn into standard -1/4x+y=-3 - Remember that in standard, all numbers must be integers, and A has to be positive, so multiply by -4 x-4y=12

OpenStudy (anonymous):

Y=m(x)+b m=slope b=y intercept

OpenStudy (anonymous):

Standard form is in the form of Ax+By=C, all numbers being integers and A being positive, I believe.

OpenStudy (anonymous):

Slope intercept: y=mx+b Point Slope: (y-y)=m(x-x) Standard: Ax+By=C

OpenStudy (anonymous):

Oh right, I didn't see the standard form part. So yes, Ax+By=C

OpenStudy (anonymous):

Thank You :)